# Diffusion Models Are Real-Time Game Engines

### v1

@@ 7 Discussion
Summary. We introduced GameNGen, and demonstrated that high-quality real-time game play at 20 frames per second is possible on a neural model. We also provided a recipe for converting an interactive piece of software such as a computer game into a neural model.
@@
Limitations. GameNGen suffers from a limited amount of memory. The model only has access to a little over 3 seconds of history, so it’s remarkable that much of the game logic is persisted for drastically longer time horizons. While some of the game state is persisted through screen pixels (e.g. ammo and health tallies, available weapons, etc.), the model likely learns strong heuristics that allow meaningful generalizations. For example, from the rendered view the model learns to infer the player’s location, and from the ammo and health tallies, the model might infer whether the player has already been through an area and defeated the enemies there. That said, it’s easy to create situations where this context length is not enough. Continuing to increase the context size with our existing architecture yields only marginal benefits (Section 5.2.1), and the model’s short context length remains an important limitation. The second important limitation are the remaining differences between the agent’s behavior and those of human players. For example, our agent, even at the end of training, still does not explore all of the game’s locations and interactions, leading to erroneous behavior in those cases.
@@
Future Work. We demonstrate GameNGen on the classic game DOOM. It would be interesting to test it on other games or more generally on other interactive software systems; We note that nothing in our technique is DOOM specific except for the reward function for the RL-agent. We plan on addressing that in a future work; While GameNGen manages to maintain game state accurately, it isn’t perfect, as per the discussion above. A more sophisticated architecture might be needed to mitigate these; GameNGen currently has a limited capability to leverage more than a minimal amount of memory. Experimenting with further expanding the memory effectively could be critical for more complex games/software; GameNGen runs at 20 or 50 FPS222Faster than the original game DOOM ran on the some of the authors’ 80386 machines at the time! on a TPUv5. It would be interesting to experiment with further optimization techniques to get it to run at higher frame rates and on consumer hardware.
@@
Towards a New Paradigm for Interactive Video Games. Today, video games are programmed by humans. GameNGen is a proof-of-concept for one part of a new paradigm where games are weights of a neural model, not lines of code. GameNGen shows that an architecture and model weights exist such that a neural model can effectively run a complex game (DOOM) interactively on existing hardware. While many important questions remain, we are hopeful that this paradigm could have important benefits. For example, the development process for video games under this new paradigm might be less costly and more accessible, whereby games could be developed and edited via textual descriptions or examples images. A small part of this vision, namely creating modifications or novel behaviors for existing games, might be achievable in the shorter term. For example, we might be able to convert a set of frames into a new playable level or create a new character just based on example images, without having to author code. Other advantages of this new paradigm include strong guarantees on frame rates and memory footprints. We have not experimented with these directions yet and much more work is required here, but we are excited to try! Hopefully this small step will someday contribute to a meaningful improvement in people’s experience with video games, or maybe even more generally, in day-to-day interactions with interactive software systems.
@@ Acknowledgements
We’d like to extend a huge thank you to Eyal Segalis, Eyal Molad, Matan Kalman, Nataniel Ruiz, Amir Hertz, Matan Cohen, Yossi Matias, Yael Pritch, Danny Lumen, Valerie Nygaard, the Theta Labs and Google Research teams, and our families for insightful feedback, ideas, suggestions, and support.
@@
Dani Valevski: Developed much of the codebase, tuned parameters and details across the system, added autoencoder fine-tuning, agent training, and distillation.
Yaniv Leviathan: Proposed project, method, and architecture, developed the initial implementation, key contributor to implementation and writing.
Moab Arar: Led auto-regressive stabilization with noise-augmentation, many of the ablations, and created the dataset of human-play data.
Shlomi Fruchter: Proposed project, method, and architecture. Project leadership, initial implementation using DOOM, main manuscript writing, evaluation metrics, random policy data pipeline.
Correspondence to shlomif@google.com and leviathan@google.com.
@@ A.2 Fine-Tuning Latent Decoder Examples
A.3 Reward Function
@@
A.4 Reducing Inference Steps
@@
| Steps | PSNR | LPIPS |
| D | ||
| 1 | ||
| 2 | ||
| 4 | ||
| 8 | ||
| 16 | ||
| 32 | ||
| 64 |
A.5 Agent vs Random Policy
@@
A.6 Human Eval Tool

### v2

@@ 7 Discussion
Summary. We introduced GameNGen, and demonstrated that high-quality real-time gameplay at 20 frames per second is possible on a neural model.
@@
Limitations. (1) GameNGen suffers from a limited amount of memory. The model only has access to a little over 3 seconds of history, so it’s remarkable that much of the game logic is persisted for drastically longer time horizons (see the supplementary for multi-minute gameplay). While some of the game state is persisted through screen pixels (e.g. ammo and health tallies, available weapons, etc.), the model likely learns strong heuristics that allow meaningful generalizations. For example, the model infers the current location from the rendered view, and can guess if enemies in an area are defeated from the ammo and health tallies. That said, it’s easy to create situations where this context length is not enough (see video in the supplementary). Also, these heuristics might be wrong, for example, if the player repeatedly shoots, GameNGen might decide to spawn an enemy, as the agent would usually shoot when enemies are present. Continuing to increase the context size with our existing architecture yields only marginal benefits (Section 5.2.1), and the model’s short context length remains an important limitation. (2) The second important limitation are the remaining differences between the agent’s behavior and those of human players. For example, our agent, even at the end of training, still does not explore all of the game’s locations and interactions, leading to erroneous behavior in those cases (see video in the supplementary). (3) Another important limitation, is that we are not able to easily produce new games with GameNGen. Like traditional game engines, GameNGen interactively runs the game-loop (update state based on input and render it to the screen). However, most game engines offer another important feature which is the ability to easily create new games, which GameNGen currently lacks.
@@
Future Work. We demonstrate GameNGen on the classic game DOOM. It would be interesting to test it on other games or more generally on other interactive software systems; We note that nothing in our technique is DOOM specific except for the reward function for the RL-agent. We plan on addressing that in a future work; While GameNGen manages to maintain game state accurately, it isn’t perfect, as per the discussion above. A more sophisticated architecture and training scheme might be needed to mitigate these; GameNGen currently has a limited capability to leverage more than a minimal amount of memory. Experimenting with further expanding the memory effectively could be critical for more complex games/software; GameNGen runs at 20 or 50 FPS222Faster than the original game DOOM ran on the some of the authors’ 80386 machines at the time! on a TPUv5. It would be interesting to experiment with further optimization techniques to get it to run at higher frame rates and on consumer hardware.
@@
Towards a New Paradigm for Interactive Video Games. Today, video games are programmed by humans. GameNGen is a proof-of-concept for one part of a new paradigm where games are weights of a neural model, not lines of code. GameNGen shows that an architecture and model weights exist such that a neural model can effectively run a complex game (DOOM) interactively on existing hardware. While many important questions remain, we are hopeful that this paradigm could have important benefits. For example, the development process for video games under this new paradigm might be less costly and more accessible, whereby games could be developed and edited via textual descriptions or examples images. A small part of this vision, namely creating modifications or novel behaviors for existing games, might be achievable in the shorter term. For example, we might be able to convert a set of frames into a new playable level or create a new character just based on example images, without having to author code (see Appendix A.4). Other advantages of this new paradigm include strong guarantees on frame rates and memory footprints. We have not experimented with these directions yet and much more work is required here, but we are excited to try! Hopefully this small step will someday contribute to a meaningful improvement in people’s experience with video games, or maybe even more generally, in day-to-day interactions with interactive software systems.
Broader Impact
Societal impact. GameNGen demonstrates that it is possible to simulate interactive games in real-time using neural networks, opening up new possibilities for game development. Similarly to other generative technologies like LLMs, text-to-image and text-to-video models, it will be important to explore how to empower users and game developers to build new experiences responsibly.
Reproducibility. We prioritized reproducibility in our implementation choices. We opt to use a relatively small open-source and open-weights foundation model (Stable Diffusion 1.4) which is widely accessible for fine-tuning. The game environment we use, VizDoom, is well-documented and open-source. Finally, we include detailed descriptions of training parameters and data generation configurations, and share performance metrics as a baseline for future work.
@@ Acknowledgements
We’d like to extend a huge thank you to Eyal Segalis, Eyal Molad, Matan Kalman, Nataniel Ruiz, Amir Hertz, Matan Cohen, Yossi Matias, Yael Pritch, Danny Lumen, Valerie Nygaard, Michelle Tadmor Ramanovich, the Theta Labs and Google Research teams, and our families for insightful feedback, ideas, suggestions, and support.
@@
Shlomi and Yaniv proposed the project, method, and architecture, developed the initial implementation, and made key contributions to implementation and writing. Dani developed much of the final codebase, tuned parameters and details across the system, added autoencoder fine-tuning, agent training, and distillation. Moab led auto-regressive stabilization with noise-augmentation, many of the ablations, and created the dataset of human-play data.
@@ A.2 Fine-Tuning Latent Decoder Examples
A.3 Dataset Size Analysis
[…]
A.4 Out-of-Distribution Sampling
To further explore GameNGen’s ability to generate behaviors not present in the training data, we performed initial experiments by taking frames from the game, editing them with a graphical picture editor, and starting the generation from the edited frames. Specifically, we replicate the same frame for the entirety of the history buffer, with the “no key pressed” action. To encourage further generalization, in this setup we train a new GameNGen model where we randomize the agent’s starting location. Due to the noise augmentation (Section 3.2.1), small local changes get ignored. When performing more substantial changes, the model usually generates unseen levels and situations. For example, Fig. 14 shows examples of pasting a game character into areas where they do not appear in the training data (e.g., inserting a monster from an advanced level into an early one). We observe that the model often consistently integrates the added characters into the new location, and they move, shoot at the player, and cause damage. Figure 15 demonstrates modifying a level’s layout by inserting features such as walls, doors, or pools from other areas. The model successfully integrates these into the environment, rendering new viewpoints as the player navigates. We hope to further explore these preliminary results in future work.
A.5 Reward Function
@@
A.6 Reducing Inference Steps
@@
A.7 Agent vs Random Policy
@@
A.8 Human Eval Tool
A.9 Additional related work
Interactive 3D Simulation. Interactive 2D and 3D simulations are well-developed in computer graphics (Akenine-Mller et al., 2018). Game engines like Unreal and Unity a stream of images based on user input, tracking world state (e.g., player position, objects, lighting) and game logic (e.g., score). While film productions uses computationally intensive ray tracing (Shirley & Morley, 2008), game engines prioritize speed (30-60 FPS) with optimized polygon rasterization via GPUs. Physical effects, such as shadows and lighting, are approximated for efficiency rather than simulated with full accuracy.
Neural 3D Simulation. Neural methods for reconstructing 3D representations have made significant advances over the last years. NeRFs (Mildenhall et al., 2020) parameterize radiance fields using a deep neural network that is specifically optimized for a given scene from a set of images taken from various camera poses. Once trained, novel point of views of the scene can be sampled using volume rendering methods. Gaussian Splatting (Kerbl et al., 2023) approaches build on NeRFs but represent scenes using 3D Gaussians and adapted rasterization methods, unlocking faster training and rendering times. While demonstrating impressive reconstruction results and real-time interactivity, these methods are often limited to static scenes.
Video Diffusion Models. Diffusion models achieved state-of-the-art results in text-to-image generation (Saharia et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Podell et al., 2023), a line of work that has also been applied for text-to-video generation tasks (Ho et al., 2022; Blattmann et al., 2023b; a; Gupta et al., 2023; Girdhar et al., 2023; Bar-Tal et al., 2024). Despite impressive advancement in realism, text adherence and temporal consistency, video diffusion models remain too slow for real-time applications. Our work extends this line of work and adapts it for real-time generation conditioned autoregressively on a history of past observations and actions.
A.10 Simulating a platform game
We experimented with the simple platform game ”Chrome Dino” to demonstrate GameNGen’s ability to simulate a different game type (see Fig. 18). For data gathering, we train an RL-Agent utilizing the Deep Q-Network (DQN Mnih et al. (2015a)) algorithm with experience replay, where the reward function is derived from the in-game score. The agent selects actions based on an epsilon greedy policy, with a decaying factor of 0.9995. During training, we recorded 2K episodes and used them to train the diffusion model using the same settings detailed in the main paper, with the following modifications: (1) we used a 32-frame context, (2) a resolution of 256x512, and (3) performed only 3,000 training steps. Similar to the results with DOOM, the simulation generated by GameNGen is fully playable over long trajectories in real-time, with visual quality comparable to the source game (see example video in the supplementary).
In addition, GameNGen supports simulating game-session management actions, such as game termination and automatic game replay. To achieve this, we concatenated two randomly selected episodes and sampled 32 frames from the combined sequence. This approach allows transitions between sessions to be represented when the sampled context spans both episodes.